OpenStudy (anonymous):

u substitution, anti derivative of (x+1)(2-x)^1/2 dx

OpenStudy (mr.math):

Put \(u=2-x\).

OpenStudy (anonymous):

i know that

OpenStudy (mr.math):

The rest is easy. \(u=2-x \implies x=2-u \implies dx=-du\), then the integral becomes \[\int (2-u+1)u^{\frac{1}{2}} (-du)=\int (-3u^{\frac{1}{2}}+u^{\frac{3}{2}})du=\cdots\]

OpenStudy (anonymous):

I know whatI did wrong, my dx was was wrong since I had du/1 when it should of been over -1